OpenStudy (anonymous):

What is a reasonable estimate for the area of a circle with radius 4 cm? A. A ≈ 12 cm2 B. A ≈ 16 cm2 C. A ≈ 24 cm2 D. A ≈ 48 cm2

OpenStudy (anonymous):

Area = (3.14)(4^2) = 3.14*16 = 48 something..

OpenStudy (swissgirl):

pi r^2 is area if i remember correctly

OpenStudy (swissgirl):

YAYYYYYY awesome i remembered smth at least

OpenStudy (anonymous):

Ha ha ha..